Sketch the probability density graph (not normalized) for a particle in a ground-state box (n = 1)
Answer #1

The non-normalized wave function of the particle in a box of length L, in ground state(n=1) is given by

( - ) ( ) y(x) = Asin = A sin

where A is the normalisation constant.

The probability density is

P(x)=|\psi(x)|^2=|A|^2 \sin^2 \Big( \frac{\pi x}{L}\Big)

The plot of P(x)\,\, vs\,\,x is shown below

Plot of ground state probability density of particle in a box P(x)